# namerefThe definition of as changed

### v1

@@ Abstract
We study the security of a quantum key distribution (QKD) protocol under the one-sided device-independent (1sDI) setting, which assumes trust in only one party’s measurement device. This approach effectively provides a balance between the experimental viability of device-dependent (DD-QKD) and the minimal trust assumptions of device-independent (DI-QKD). An analytical lower bound on the asymptotic key rate is derived to provide security against collective attacks, in which the eavesdropper’s information is limited only by the function of observed violation of a linear quantum steering inequality, specifically the three-setting Cavalcanti–Jones–Wiseman–Reid (CJWR) inequality. We provide a closed-form key rate formula by reducing the security analysis to mixtures of Bell-diagonal states by utilizing symmetries of the steering functional. We show that the protocol tolerates higher quantum bit error rates (QBER) than present DI-QKD protocols by benchmarking its performance under depolarizing noise. Furthermore, we explore the impact of detection inefficiencies and show that, in contrast to DI-QKD which requires near-perfect detection, secure key generation can be achieved even with lower detection efficiency on the untrusted side. These findings demonstrate the viability of using 1sDI-QKD with current technology and highlight its advantages as a steering-based substitute for secure quantum communication.
@@ I Introduction
Under ideal collective attacks, control over just one measurement device and the source is sufficient to compromise security, as demonstrated by the security analyses in Refs. [16, 27]. The 1sDI-QKD framework, which fits in between DI-QKD and SDI/MDI/DD-QKD in the hierarchy of trust models, is based on this observation and assumes trust in only one party’s device, usually Bob. The structure of 1sDI-QKD is consistent with the concept of quantum steering, which was first proposed by Schrödinger [37], and formalized later by Wiseman et al. [38]. Various criteria, such as Reid’s uncertainty-based test [39], the CJWR inequality [40, 41], entropic [42], fine-grained [43], and sum-uncertainty-relation-based formulations [44], can be used to identify steering, which captures the ability of a trusted party to nonlocally affect the state of an untrusted party. Its distinction from entanglement and Bell nonlocality has been established both theoretically [45, 46] and experimentally [40, 47], and further quantified using dedicated measures [48, 49, 50, 51]. Steering has also been shown to play an important role in certification of quantum states and measurements in the one-sided device-independent scenario [52, 53, 54].
@@ II CJWR-based 1sDI-QKD Protocol
Entanglement-based QKD provides inherent security based on quantum mechanics. The BBM92 scheme [7] depends on strong measurement correlations without utilizing nonlocality, whereas the Ekert91 protocol [4] certifies key security through violations of Bell inequality. Due to its sensitivity to loss and detection inefficiencies, DI-QKD [16, 27, 10, 25] undermines trust in all devices but is experimentally demanding. The 1sDI-QKD [55, 56], where only one party’s device is trusted (typically Bob’s), provides a practical alternative. Here, quantum steering [37, 38], an intermediate form of nonclassicality, enables security certification against an untrusted device, making 1sDI-QKD more tolerant to experimental imperfections.
@@ III Security Analysis Against Collective Attacks
We assume that Alice, Bob, and Eve share a tripartite pure state |$\ket{\Phi_{\text{ABE}}}\in\mathcal{H}_{\text{A}}^{\otimes N}\otimes\mathcal{H}_{\text{B}}^{\otimes N}\otimes\mathcal{H}_{\text{E}},$, where $N$ is the number of rounds used for key generation. Without loss of generality, we take the local Hilbert spaces to be of equal finite dimension, i.e., $\mathcal{H}_{\text{A}}\simeq\mathcal{H}_{\text{B}}\simeq\mathbb{C}^{d}$.
For the security analysis, we assume that Eve performs a collective, or i.i.d., attack [61, 62, 19, 16, 27, 63, 64, 18, 21, 65, 66, 67, 10]. In this setting, the state and measurement procedure used by Eve are the same in every round, and independent across rounds. As a result, the total shared state between Alice, Bob, and Eve takes the form |$\ket{\Phi_{\text{ABE}}}=\ket{\phi_{\text{ABE}}}^{\otimes N},$ where |$\ket{\phi_{\text{ABE}}}$ the state is shared in a single round of the protocol. In addition, we assume that the devices are memoryless, meaning the measurement in each round depends only on the current input and not on previous rounds.
@@
Step 1: To compute a tight upper bound on Eve’s accessible information, we simplify the analysis using techniques similar in spirit to those introduced in Ref. [27]. In the one-sided device-independent scenario, Bob’s measurement device is fully trusted and assumed to perform Pauli measurements. Therefore, even if Eve prepares a general state in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$, Bob effectively accesses only a two-dimensional Hilbert space. As a result, without loss of generality, we can assume that the shared state distributed by Eve is a qudit–qubit state. Certain symmetries inherent in the CJWR inequality can be used to simplify the analysis by reducing the effective dimension of the shared state. This is formalized in the following lemma.
@@
If the observables mutually anticommute, i.e., $\{A_{l},A_{m}\}=0$ for $l\neq m$, then $[A_{l},A_{m}]=2A_{l}A_{m}$. Hence,
@@
Our proof strategy mirrors the dimensionality reduction argument employed in the CHSH scenario [68], where any two dichotomic observables with eigenvalues ±$\pm 1$ are shown to generate a two-dimensional invariant subspace. In our case, the set of three mutually anticommuting observables $A_{1},A_{2},A_{3}$ similarly generate a representation of the real Clifford algebra $\mathrm{Cl}_{3}(\mathbb{R})$, whose minimal irreducible representation is two-dimensional [69]. This justifies the restriction to a two-qubit system without loss of generality. Hence, there exists a subspace $\mathcal{H}_{2}\subseteq\mathcal{H}_{A}$ such that $A_{l}$ act as $\sigma_{l}$ on $\mathcal{H}_{2}\cong\mathbb{C}^{2}$.
@@
Step 2: In the previous step, we argued that, without loss of generality, Eve can restrict herself to preparing a bipartite state in $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, where Alice’s measurements are fixed to be qubit Pauli observables. We now investigate which class of $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ states enables Eve to extract the maximum possible information, while maintaining an optimal value of the CJWR expression between Alice and Bob. In the following lemma, we show that any such two-qubit state $\rho\in\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ can be reduced to a Bell-diagonal form without affecting the CJWR value.
@@
Step 3: Without loss of generality, Eve can send any mixture of Bell-diagonal state like $\rho_{AB}=\sum_{\Lambda}p_{\Lambda}\rho_{\Lambda}$, where $\Lambda$ is a classical ancilla known to her. Now, we need to calculate the Holevo bound $\chi_{\Lambda}(B_{3}:E)$ for that Bell-diagonal state. For the Bell diagonal state the Holevo bound $\chi_{\Lambda}(B_{3}:E)$ can be calculated as,
@@
with equality in the first bound if and only if either $\Lambda_{1,3}=0$ or $\Lambda_{2,4}=0$.
@@
$\mathcal{F}_{3}^{\Lambda}=\sqrt{(T^{\Lambda}_{11})^{2}+(T^{\Lambda}_{22})^{2}+(T^{\Lambda}_{33})^{2}}.$
As shown in Step 3 under optimal attack, this can be compactly written as $\mathcal{F}_{3}^{\Lambda}=\sqrt{4\mathcal{R}^{2}-1}$, where $\mathcal{R}$ is defined in terms of the Bell-state probabilities and the threshold $\mathcal{R}=1/\sqrt{2}$ corresponds to the classical limit for the CJWR inequality, analogous to the CHSH case in Ref. [27].
@@ IV Noise tolerance of CJWR-based 1sDI-QKD
To quantitatively compare the noise tolerance of 1sDI-QKD against DI and DD protocols, we consider a widely used noise model in the QKD literature [10, 70, 71, 72, 16, 27, 55]. In this model, the maximally entangled Bell state |$\ket{\phi^{+}}$ is subjected to depolarizing noise, resulting in the mixed state
$\rho_{\nu}=\nu\ket{\phi^{+}}\bra{\phi^{+}}+(1-\nu)\frac{\mathbb{I}}{4},$ (14)
@@ V Detection Efficiency in 1sDI-QKD
In the 1sDI scenario, detection inefficiency manifests through no-click events on Alice’s side, which we denote by the null outcome $\varnothing$. This effectively increases Alice’s output alphabet to three possible outcomes: $\{+1,-1,\varnothing\}$. Following the approach of Ref. [27], we address this by deterministically mapping the null outcome to $-1$, thereby reducing the measurement to a binary-output POVM. The resulting effective measurement operators on Alice’s side take the form
@@
where $\eta_{A}\in[0,1]$ denotes Alice’s detection efficiency, and $A_{\pm 1|i}$ are the ideal POVM elements for input $i$.
@@
whereas under postselection, the key rate is given by
@@
To explore how visibility impacts the security threshold, we further examine the relationship between $\nu$ and the minimum detection efficiency needed for key generation. As shown in Fig. 4, the threshold $\eta_{A}$ decreases with increasing visibility ($\nu$). For all values of $\nu$, the post-selected strategy performs better, consistently allowing secure key generation at lower detection efficiencies. At $\nu=1$, we recover the earlier thresholds from Fig. 3, confirming consistency between the two analyses.
When compared with other known QKD protocols, our approach remains competitive. In entanglement-based protocols with a trusted Bob, critical detection efficiencies are around $89.6\%$ without postselection and $83.3\%$ with postselection [55, 18]. In the DI setting under collective attacks, the requirements are even stricter $92.3\%$ without postselection and $88.9\%$ with it [27]. The original one sided DI-QKD protocol proposed by Branciard et al. achieves lower thresholds of $78\%$ (without postselection) and $65.9\%$ (with postselection) using a two-setting BBM92 like scheme and an entropic uncertainty-based proof [55]. Under ideal detection conditions, the key rate obtained from this protocol matches that of the standard entanglement-based BB84 or BBM92 protocols analyzed under coherent attacks [17, 27, 10].
@@
Our protocol is inspired by the Ekert91 framework and is based on the violation of the CJWR steering inequality [40]. The security is established not via Bell inequality violation, but through quantum steering, which provides a robust and realistic alternative when only one device is trusted. The key rate is analytically derived assuming collective attacks, and the bound on Eve’s information is obtained following the approach in Ref. [16]. Notably, the critical detection efficiency in steering based protocols depends on the chosen steering inequality. While we use a fixed bound, experimental studies suggest that the threshold can be adapted based on observed efficiencies [59], potentially pushing the security threshold even lower.

### v2

# Secure One-Sided Device-Independent Quantum Key Distribution Under Collective Attacks with Enhanced Robustness

@@ Abstract
We study the security of a quantum key distribution (QKD) protocol under the one-sided device-independent (1sDI) setting, which assumes trust in only one party’s measurement device. This approach effectively provides a balance between the experimental viability of device-dependent (DD-QKD) and the minimal trust assumptions of device-independent (DI-QKD). An analytical lower bound on the asymptotic key rate is derived to provide security against collective attacks, in which the eavesdropper’s information is limited only by the function of observed violation of a linear quantum steering inequality, specifically the three-setting Cavalcanti–Jones–Wiseman–Reid (CJWR) inequality. We provide a closed-form key rate formula by reducing the security analysis to mixtures of Bell-diagonal states by utilizing symmetries of the steering functional. We show that the protocol tolerates higher quantum bit error rates (QBER) than present DI-QKD protocols by benchmarking its performance under depolarizing noise. Furthermore, we explore the impact of detection inefficiencies and show that, in contrast to DI-QKD, which requires near-perfect detection, secure key generation can be achieved even with lower detection efficiency on the untrusted side. These findings highlight the advantages of 1sDI-QKD as a steering-based alternative for secure quantum communication and provide insights relevant for near-future experimental implementations.
@@ I Introduction
Under ideal collective attacks, control over just one measurement device and the source is sufficient to compromise security, as demonstrated by the security analyses in Refs. [16, 27]. The 1sDI-QKD framework, which fits in between DI-QKD and SDI/MDI/DD-QKD in the hierarchy of trust models, is based on this observation and assumes trust in only one party’s device, usually Bob. The structure of 1sDI-QKD is consistent with the concept of quantum steering, which was first proposed by Schrödinger [37], and formalized later by Wiseman et al. [38]. Various criteria, such as Reid’s uncertainty-based test [39], the CJWR inequality [40, 41], entropic [42], fine-grained [43], and sum-uncertainty-relation-based formulations [44], can be used to identify steering, which captures the ability of a trusted party to nonlocally affect the state of an untrusted party. Its distinction from entanglement and Bell nonlocality has been established both theoretically [45, 46] and experimentally [40, 47], and further quantified using dedicated measures [48, 49, 50, 51].
@@ II CJWR-based 1sDI-QKD Protocol
Entanglement-based QKD provides inherent security based on quantum mechanics. The BBM92 scheme [7] depends on strong measurement correlations without utilizing nonlocality, whereas the Ekert91 protocol [4] certifies key security through violations of a Bell inequality. Due to its sensitivity to loss and detection inefficiencies, DI-QKD [16, 27, 10, 25] undermines trust in all devices but is experimentally demanding. The 1sDI-QKD [52, 53], where only one party’s device is trusted (typically Bob’s), provides a practical alternative. Here, quantum steering [37, 38], an intermediate form of nonclassicality, enables security certification against an untrusted device, making 1sDI-QKD more tolerant to experimental imperfections.
@@ III Security Analysis Against Collective Attacks
[…]
@@
Step 1: To compute a tight upper bound on Eve’s accessible information, we simplify the analysis using techniques similar in spirit to those introduced in Ref. [27]. In the one-sided device-independent scenario, Bob’s measurement device is fully trusted and assumed to perform Pauli measurements. Therefore, even if Eve prepares a general state in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$, Bob effectively accesses only a two-dimensional Hilbert space. As a result, without loss of generality, we can assume that the shared state distributed by Eve is a qudit-qubit state. Certain symmetries inherent in the CJWR inequality can be used to simplify the analysis by reducing the effective dimension of the shared state. This is formalized in the following lemma.
@@
If the observables mutually anticommute, i.e., $\{A_{l},A_{m}\}=0$ for $l\neq m$, then $A_{l},A_{m}]=2A_{l}A_{m}$. Hence,
@@
Our proof strategy mirrors the dimensionality reduction argument employed in the CHSH scenario [65], where any two dichotomic observables with eigenvalues $\pm 1$ are shown to generate a two-dimensional invariant subspace. In our case, the set of three mutually anticommuting observables $A_{1},A_{2},A_{3}$ similarly generate a representation of the real Clifford algebra $\mathrm{Cl}_{3}(\mathbb{R})$, whose minimal irreducible representation is two-dimensional [66]. This justifies the restriction to a two-qubit system without loss of generality. Hence, there exists a subspace $\mathcal{H}_{2}\subseteq\mathcal{H}_{A}$ such that $A_{l}$ act as $\sigma_{l}$ on $\mathcal{H}_{2}\cong\mathbb{C}^{2}$.
@@
Step 2: In the previous step, we argued that, without loss of generality, Eve can restrict herself to preparing a bipartite state in $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, where Alice’s measurements are fixed to be qubit Pauli observables. We now investigate which class of $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ states enables Eve to extract the maximum possible information while maintaining an optimal value of the CJWR expression between Alice and Bob. In the following lemma, we show that any such two-qubit state $\rho\in\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ can be reduced to a Bell-diagonal form without affecting the CJWR value.
@@
Step 3: Without loss of generality, Eve can send any mixture of Bell-diagonal states like $\rho_{AB}=\sum_{\Lambda}p_{\Lambda}\rho_{\Lambda}$, where $\Lambda$ is a classical ancilla known to her. Now, we need to calculate the Holevo bound $\chi_{\Lambda}(B_{3}:E)$ for that Bell-diagonal state. For the Bell diagonal state the Holevo bound $\chi_{\Lambda}(B_{3}:E)$ can be calculated as,
@@
with equality in the first bound if and only if either $\Lambda_{1,3}=0$ or $\Lambda_{2,4}=0$ [16].
@@
$\mathcal{F}_{3}^{\Lambda}=\sqrt{(T^{\Lambda}_{11})^{2}+(T^{\Lambda}_{22})^{2}+(T^{\Lambda}_{33})^{2}}.$ (12)
For the optimal collective attack, the entropic bound of Ref. [16], as mentioned in step 3, is saturated by extremal Bell-diagonal states satisfying either $\Lambda_{1}=\Lambda_{3}=0$ or $\Lambda_{2}=\Lambda_{4}=0$. Without loss of generality, we consider the first case and write
$\Lambda_{1}$ $=\Lambda_{3}=0,\qquad\Lambda_{2}=k,\qquad\Lambda_{4}=1-k,$
with $k\in[0,1]$. Then the parameter $R$ entering the entropic bound is then
$R^{2}$ $=(\Lambda_{1}-\Lambda_{2})^{2}+(\Lambda_{3}-\Lambda_{4})^{2}$
$=k^{2}+(1-k)^{2}.$
Using the correlation matrix elements of a Bell-diagonal state, one finds
$T_{11}^{\Lambda}$ $=1-2k,\qquad T_{22}^{\Lambda}=1-2k,\qquad T_{33}^{\Lambda}=-1,$
so that the CJWR steering value of Eq. (12) becomes
$\mathcal{(}F_{3}^{\Lambda})^{2}$ $=2(1-2k)^{2}+1$
$=8k^{2}-8k+3$
$=4(k^{2}+(1-k)^{2})-1.$
On the other hand,
$4R^{2}-1$ $=4\big[k^{2}+(1-k)^{2}\big]-1$
$=8k^{2}-8k+3,$
which yields
$\mathcal{(}F_{3}^{\Lambda})^{2}=4R^{2}-1.$
The complementary extremal case $\Lambda_{2}=\Lambda_{4}=0$ can be treated analogously and leads to the same relation. Hence,
$R=\frac{1}{2}\sqrt{1+\mathcal{(}F_{3}^{\Lambda})^{2}},$
demonstrating that the intermediate parameter $R$ is fully determined by the observed CJWR steering violation $\mathcal{F}_{3}^{\Lambda}$.
This can be compactly written as $\mathcal{F}_{3}^{\Lambda}=\sqrt{4\mathcal{R}^{2}-1}$, where $\mathcal{R}$ is defined in terms of the Bell-state probabilities and the threshold $\mathcal{R}=1/\sqrt{2}$ corresponds to the classical limit for the CJWR inequality, analogous to the CHSH case in Ref. [27].
@@ IV Noise tolerance of CJWR-based 1sDI-QKD
[…]
@@
Reference Protocol type Attack model Role of steering Key features
Branciard et al. [52] Prepare-measure (BBM92-type), 1sDI Coherent (memoryless) Implicit: steering required for security when the protocol is expressed in an entanglement-based picture Eve’s information bounded via QBERs in two bases using entropic uncertainty relations; reduces to standard DD key rate under symmetric noise
Tomamichel et al. [53] Prepare-measure (BB84-type), 1sDI Coherent, finite-size Implicit: steering arises through monogamy-of-entanglement games Fully composable finite-size security; very low key rates, positive only at very low QBER
Pramanik et al. [43] Entanglement-based, 1sDI Individual attacks Explicit: violation of a steering inequality guarantees security Provides a lower bound on the key rate from steering violation; no explicit Holevo-information bound
Mukherjee et al. [54] Entanglement-based, 1sDI Asymptotic Explicit: CJWR steering violation combined with local filtering Security is characterized via an effective QBER after filtering; Eve’s information not bounded directly
Masini & Sarkar [55] Prepare–measure (BB84-type), 1sDI Coherent, finite-size Implicit: steering used to bound Eve’s entropy via SDP and EAT Fully composable; numerical bounds only, no closed-form analytic dependence on QBER or steering
This work Entanglement-based (E91-type), 1sDI Collective Explicit: security certified directly by CJWR steering violation The analytic Holevo bound expressed in terms of the observed pair $(Q,\mathcal{F}_{3})$; model-independent detection-efficiency thresholds
Table 1: Comparison of representative 1sDI-QKD protocols, highlighting protocol type, attack model, and the operational role of steering.
Our protocol differs conceptually from most existing 1sDI-QKD schemes summarized in Table 1. Whereas earlier approaches ultimately bound Eve’s information using one or more QBER parameters, our analysis ties security directly to the observed strength of steering correlations rather than relying on error rates alone. In particular, in the protocol of Branciard et al. [52], Eve’s information is bounded via an entropic uncertainty relation using QBERs measured in two complementary bases; under ideal detection efficiency and symmetric depolarizing noise, these QBERs coincide, causing the 1sDI key-rate bound to reduce to the standard device-dependent BB84/BBM92 expression [17, 27, 10] r=12h(Q),r=1-2h(Q), with a critical QBER $Q_{c}\simeq 11\%$. A similar QBER-centric behavior appears in the finite-size, monogamy-game-based analysis of Tomamichel et al. [53], where a positive key can be obtained only for very low depolarizing noise, $Q\lesssim 1.5\%$. Steering-based QKD protocols that do not explicitly employ a Holevo-information bound, notably those of Pramanik et al. [43] (restricted to individual attacks) and Mukherjee et al. [54] (with no adversarial model specified), certify security through the violation of a steering inequality. In these works, the condition for a positive key rate can be re-expressed in terms of an effective QBER, leading to a comparatively high tolerable noise threshold, as $Q_{c}\simeq 21\%$. Nevertheless, since Eve’s information is not explicitly bounded, this threshold does not correspond to a quantitative characterization of Eve’s accessible information. While Masini et al. [55] mainly focused on demonstrating near-optimal detection-efficiency thresholds by means of SDP-based entropy bounds and a three-outcome measurement model, their approach yields purely numerical key-rate bounds without a closed-form dependence on QBER or steering violation. In contrast, our E91-style 1sDI protocol bounds Eve’s Holevo information directly from the observed pair $(Q,\mathcal{F}_{3})$, leading under the same depolarizing-noise model to a nontrivial critical QBER $Q_{c}\simeq 8.62\%$ derived solely from measured correlations and without assuming a noise model and measurement model.
@@ V Detection Efficiency in 1sDI-QKD
In the 1sDI scenario, detection inefficiency manifests through no-click events on Alice’s side, which we denote by the null outcome $\varnothing$. This effectively increases Alice’s output alphabet to three possible outcomes: $\{+1,-1,\varnothing\}$. Following the approach of Ref. [27], we address this by deterministically mapping the null outcome to 1-1, thereby reducing the measurement to a binary-output POVM. The resulting effective measurement operators on Alice’s side take the form
@@
where $\eta_{A}\in[0,1]$ denotes Alice’s detection efficiency and $A_{\pm 1|i}$ are the ideal POVM elements for input $i$.
@@
whereas under post-selection, the key rate is given by
@@
To explore how visibility impacts the security threshold, we further examine the relationship between $\nu$ and the minimum detection efficiency needed for key generation. As shown in Fig. 4, the threshold $\eta_{A}$ decreases with increasing visibility ($\nu$). For all values of $\nu$, the post-selected strategy performs better, consistently allowing secure key generation at lower detection efficiencies. At $\nu=1$, we recover the earlier thresholds from Fig. 3, confirming consistency between the two analyses. We emphasize that the analysis extends beyond the ideal-visibility limit. Figure 4 shows that the detection-efficiency requirements become increasingly stringent as the visibility decreases in the experimentally relevant nonideal regime $\nu<1$. For example, at $\nu\approx 0.95$ the minimum detection efficiency is approximately $\eta_{A}\approx 78\%$ with post-selection and $\eta_{A}\approx 86\%$ without post-selection, increasing to $\eta_{A}\approx 82\%$ and $\eta_{A}\approx 90\%$, respectively, at $\nu\approx 0.90$. Consistently, Fig. 3 indicates that a positive secret key rate is achievable only for $\eta_{A}\gtrsim 74.5\%$ with post-selection and $\eta_{A}\gtrsim 82.7\%$ without post-selection, corresponding to the high-visibility regime. These results highlight the experimentally relevant trade-off between source visibility and detector efficiency in non-ideal operating conditions.
When compared with other known QKD protocols, our approach remains competitive. In entanglement-based protocols with a trusted Bob, critical detection efficiencies are around $89.6\%$ without postselection and $83.3\%$ with postselection [52, 18]. In the DI setting under collective attacks, the requirements are even stricter $92.3\%$ without postselection and $88.9\%$ with it [27]. The original one-sided DI-QKD protocol proposed by Branciard et al. achieves lower thresholds of $78\%$ (without postselection) and $65.9\%$ (with postselection) using a two-setting BBM92-like scheme and an entropic uncertainty-based proof [52]. Masini et al. [55] further reduce the required detection efficiency on the untrusted side to 50.1% by retaining nondetection events as an explicit third outcome and bounding Eve’s information numerically using SDP-based entropy-accumulation techniques. Although this three-outcome treatment avoids post-selection and enables security against coherent attacks, it significantly weakens the effective correlations, leading to reduced key rates near the threshold and security bounds that are purely numerical.
@@
Notably, the critical detection efficiency in steering-based protocols depends on the chosen steering inequality. While we use a fixed bound, experimental studies suggest that the threshold can be adapted based on observed efficiencies [56], potentially pushing the security threshold even lower.